CAN SOMEONE PLEASE HELP ME AND SHOW ME HOW TO WORK OUT THESE PROBLEMS.I IF U BREAK IT DOWN FOR ME I WILL UNDERSTAND IT BETTER.
1) 1/4+1/6-1/36=

2) WHAT VALUES OF X SOLVE
X(2 IS EXPONENT)-5X+6=0?

3) SOLVE THE FOLLOWING SIMULTANEOUS EQUATIONS
4X+2Y=28
3X-Y=11


Found 2 solutions by cyeannakis, stanbon:
Answer by cyeannakis(1)

1/4+1/6-1/36
Find the common denominator...it's 36. 36 can be divided by 4 and 6 with no remainder.
Now, give all fractions a denominator of 36 by multiplying by fraction names for one...
1/4(9/9) +1/6(6/6) - 1/36. Multiply by the fraction names for one
9/36+6/36-1/36. Now add and subtract.

Answer by stanbon(75887)

1) 1/4+1/6-1/36= x
Multiply thru by 36 to get:
9 + 6 - 1 = 36x
14 = 36x
x = 14/36 = 7/18
=======================
2) WHAT VALUES OF X SOLVE
X(2 IS EXPONENT)-5X+6=0?
x^2 - 5x + 6 = 0
Factor:
(x-3)(x-2) = 0
x = 3 or x = 2
=======================
3) SOLVE THE FOLLOWING SIMULTANEOUS EQUATIONS
4X+2Y=28
3X-Y=11
----
Divide thru the top equation by 2:
2x + y = 14
3x - y = 11
----
Add and solve for "x":
5x = 25
x = 5
----
Solve for "y":
2x+y = 14
2*5 + y = 14
y = 4
================
Cheers,
Stan H.
================